OpenStudy (anonymous):

simplify... square root 28, times, square root 63

OpenStudy (saifoo.khan):

42

OpenStudy (anonymous):

How did you do that? Did you just multiply 28 by 63 and then find the square root of the answer?

OpenStudy (mathmate):

sqrt(28)*sqrt(63) =sqrt(28*63) =sqrt(4^2*7^2*3^2) =4*7*3 =42